# Hidden yet quantifiable: A lower bound for confounding strength using randomized trials

### v2

@@ A.1 Remaining proofs
$\lim_{n_{\operatorname{rct}}\to\infty}P_{H_{0}}\left(\hat{T}_{\Gamma}^{+}\leq z_{\alpha/2}\right)$ $\leq\lim_{n_{\operatorname{rct}}\to\infty}P_{H_{0}}\left(\frac{\sqrt{n_{\operatorname{rct}}}(\hat{\mu}^{+}_{\Gamma}-\hat{\mu})-\mu_{\Gamma}^{+}\left(\mathbb{P}^{\operatorname{rct}}_{X}\right)+\mu\left(\mathbb{P}^{\operatorname{rct}}_{X},\mathbb{P}^{\operatorname{rct}}_{\operatorname{full}}\right)}{\sqrt{(\sigma_{\Gamma}^{+})^{2}+\sigma^{2}-2\xi_{\Gamma}}}\leq z_{\alpha/2}\right)$
@@
$\lim_{n\to\infty}P_{H_{0}}\left(\hat{T}_{\Gamma}^{+}\leq z_{\alpha/2}\right)$ $=\lim_{n\to\infty}P_{H_{0}}\left(\frac{\hat{\mu}^{+}_{\Gamma}-\hat{\mu}}{\sqrt{(\hat{\sigma}^{+}_{\Gamma})^{2}+\hat{\sigma}^{2}}}\leq z_{\alpha/2}\right)$
$\leq\lim_{n\to\infty}P_{H_{0}}\left(\frac{\hat{\mu}^{+}_{\Gamma}-\hat{\mu}-\mu_{\Gamma}^{+}\left(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X}\right)+\mu\left(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X},\mathbb{P}^{\operatorname{os}}_{\operatorname{full}}\right)}{\sqrt{(\hat{\sigma}^{+}_{\Gamma})^{2}+\hat{\sigma}^{2}}}\leq z_{\alpha/2}\right)$
@@ A.3 Limitations of MSM
We discuss here the intuition behind the tightness of $\hat{\Gamma}_{\text{LB}}$ in the infinite-sample limit, though it carries over to finite samples.
@@
$\Gamma_{\text{LB}}:=\inf\left\{\Gamma:\lim_{n\to\infty}\hat{\phi}_{\widetilde{\mathrm{os}}}(\Gamma,\alpha)=0\right\},$
@@
$\Gamma_{\text{LB}}$ $=\inf\left\{\Gamma:\mu(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X},\mathbb{P}^{\operatorname{os}}_{\operatorname{full}})\in[\mu_{\Gamma}^{+}(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X}),\mu_{\Gamma}^{-}(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X})]\right\}$
@@
The looseness of the lower bound can be characterized by $\Delta:=\Gamma^{\star}-\Gamma_{\text{LB}}$: if $\Delta>0$ even in the infinite-sample limit, the lower bound will not be tight. We discuss two interesting cases:

### v3

@@ A.1 Remaining proofs
$\lim_{n_{\operatorname{rct}},n_{\operatorname{os}}\to\infty}\mathbb{P}_{H_{0}(\Gamma)}\left(\hat{T}_{\Gamma}^{+}\leq z_{\alpha/2}\right)$ $\leq\lim_{n_{\operatorname{rct}},n_{\operatorname{os}}\to\infty}\mathbb{P}_{H_{0}(\Gamma)}\left(\frac{\sqrt{n_{\operatorname{rct}}}(\hat{\mu}^{+}_{\Gamma}-\hat{\mu})-\mu_{\Gamma}^{+}\left(\mathbb{P}^{\operatorname{rct}}_{X}\right)+\mu\left(\mathbb{P}^{\operatorname{rct}}_{X},\mathbb{P}^{\operatorname{rct}}_{\operatorname{full}}\right)}{\sqrt{(\sigma_{\Gamma}^{+})^{2}+\sigma^{2}-2\xi_{\Gamma}}}\leq z_{\alpha/2}\right)$
@@
$\lim_{n\to\infty}\mathbb{P}_{H_{0}(\Gamma)}\left(\hat{T}_{\Gamma}^{+}\leq z_{\alpha/2}\right)$ $=\lim_{n\to\infty}\mathbb{P}_{H_{0}(\Gamma)}\left(\frac{\hat{\mu}^{+}_{\Gamma}-\hat{\mu}}{\sqrt{(\hat{\sigma}^{+}_{\Gamma})^{2}+\hat{\sigma}^{2}}}\leq z_{\alpha/2}\right)$
$\leq\lim_{n\to\infty}\mathbb{P}_{H_{0}(\Gamma)}\left(\frac{\hat{\mu}^{+}_{\Gamma}-\hat{\mu}-\mu_{\Gamma}^{+}\left(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X}\right)+\mu\left(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X},\mathbb{P}^{\operatorname{os}}_{\operatorname{full}}\right)}{\sqrt{(\hat{\sigma}^{+}_{\Gamma})^{2}+\hat{\sigma}^{2}}}\leq z_{\alpha/2}\right)$
@@ A.3 Limitations of MSM
We discuss here the intuition behind the tightness of $\hat{\Gamma}_{\texttt{LB}}$ in the infinite-sample limit, though it carries over to finite samples.
@@
$\Gamma_{\texttt{LB}}:=\inf\left\{\Gamma:\lim_{n\to\infty}\hat{\phi}_{\widetilde{\mathrm{os}}}(\Gamma,\alpha)=0\right\},$
@@
$\Gamma_{\texttt{LB}}$ $=\inf\left\{\Gamma:\mu(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X},\mathbb{P}^{\operatorname{os}}_{\operatorname{full}})\in[\mu_{\Gamma}^{+}(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X}),\mu_{\Gamma}^{-}(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X})]\right\}$
@@
The looseness of the lower bound can be characterized by $\Delta:=\Gamma^{\star}-\Gamma_{\texttt{LB}}$: if $\Delta>0$ even in the infinite-sample limit, the lower bound will not be tight. We discuss two interesting cases: